# GraphLand: Evaluating Graph Machine Learning Models on Diverse Industrial Data

### v3

@@ 5 Experiments
5.1 Models
Graph-agnostic baselines
The tasks that we consider in our benchmark, such as fraud detection and CTR prediction, are frequently tackled in industrial settings without considering the graph structure. Thus, as baselines, we use several models that do not have access to the graph structure. We call such models graph-agnostic. Comparing results between graph-agnostic and graph-aware models might reveal how much benefit using the graph structure brings for the considered task. First, as a simple baseline, we use ResMLP — an MLP augmented with skip-connections (He et al., 2016) and layer normalization (Ba et al., 2016). It has been shown that models of this type can serve as very strong baselines for industrial data with mixed numerical and categorical features (Gorishniy et al., 2021). Further, we consider GBDT models (Friedman, 2001). These models are very popular in industrial applications and are often considered to be better than neural networks at dealing with numerical features (Gorishniy et al., 2021, 2022) and regression tasks (Zhang et al., 2023), which makes them very relevant baselines for our benchmark. We use the three most popular GBDT implementations: XGBoost (Chen & Guestrin, 2016), LightGBM (Ke et al., 2017), and CatBoost (Prokhorenkova et al., 2018).
Further, we attempt to make originally graph-agnostic models stronger on our benchmark by providing them with some graph information through feature augmentation. Specifically, for each node in the graph, we aggregate features from all its one-hop neighbors in the graph and compute the mean, maximum, and minimum values of each feature, appending these statistics to the node’s original features. This mimics one step of GNN spatial graph convolution and allows graph-agnostic models to use some of the graph information that GNNs use. We call this feature augmentation strategy neighborhood feature aggregation (NFA) and denote models that use it with the -NFA suffix. A more detailed description of NFA is provided in Appendix B.
Graph neural networks
As our main models, we use four representative GNN architectures: two classic GNNs — GCN (Kipf & Welling, 2017) and GraphSAGE (Hamilton et al., 2017), and two attention-augmented GNNs — GAT (Veličković et al., 2018) and neighborhood-attention Graph Transformer (GT) (Shi et al., 2021). For all GNNs, we use the modifications from Platonov et al. (2023b) that add skip connections, layer normalization, and MLP blocks to the models. We found that these modifications often significantly improve performance on our datasets.
It has been argued recently that in some GML datasets the graph structure does not actually help solve the considered tasks (Errica et al., 2020; Li et al., 2024b; Coupette et al., 2025; Bechler-Speicher et al., 2025). Thus, when new graph datasets are introduced, it is important to experimentally verify that the graph structure in them is actually helpful. For this, the performance of graph-agnostic and graph-aware models can be compared. To avoid other factors influencing the comparison, the two models being compared should be otherwise identical except that one has access to the graph structure and the other does not. Our set of models provides two such comparisons. First, graph-agnostic models can be compared with their NFA-augmented versions which have access to graph neighborhood feature information. Second, ResMLP can be compared to GNNs: our GNNs have the exact same backbone that our ResMLP baseline has, except GNNs have spatial graph convolution modules added, which makes the comparison fair.
Graph foundation models
Recently, there has been a lot of interest in developing graph foundation models — models that can be applied to diverse graph datasets without or with minimal fine-tuning. However, we found that the current GFMs predominantly focus on text-attributed graphs and overlook the challenge of transferring a single model to datasets with different node feature sets. This prevents such models from being truly general, as graphs in real-world applications often have many different numerical and/or categorical node attributes that are important for solving the relevant tasks. We were able to find only five GFMs that support node property prediction in graphs with arbitrary node attributes: SAMGPT (Yu et al., 2025), GCOPE (Zhao et al., 2024), OpenGraph (Xia et al., 2024), GraphFM (Lachi et al., 2024), AnyGraph (Xia & Huang, 2024). Of these models, only OpenGraph and AnyGraph have publicly available weights. Thus, we evaluate these two models. We commend the researchers who make their GFMs openly available and encourage others to do the same. For both models, we use the in-context learning (ICL) setting recommended by the authors for node classification, in which the task is cast as predicting links to virtual class nodes. Further, we were able to reproduce the pretraining of GCOPE, and we evaluate this model using the fine-tuning (FT) setting recommended by the authors. However, neither of the models supports node regression; thus, we are only able to evaluate these models on our node classification datasets. Further, we found that these GFMs cannot scale to large datasets.
5.2 Low label rate, random data split, transductive setting
@@
Results for experiments under the random low label rate data split (RL) are provided in Table 1.
@@
First, the results demonstrate that the graph structure in our datasets is very beneficial for the considered tasks as all GNNs always significantly outperform ResMLP and NFA-augmented models always significantly outperform their counterparts without NFA. Providing graph-agnostic models with graph information through NFA significantly improves their performance and even allows LightGBM-NFA to achieve the best results on four datasets including three regression ones. At the same time, GNNs outperform ResMLP-NFA on these datasets, which suggests that the success of LightGBM-NFA is due to GBDTs being better at handling numerical features and regression targets (Gorishniy et al., 2021, 2022; Zhang et al., 2023). This highlights that GBDTs with augmented features are strong baselines in realistic industrial settings. Still, GNNs manage to outperform them on most datasets, showing their effectiveness for industrial applications.
@@
As for GFMs, we find that OpenGraph, AnyGraph, and GCOPE fail to achieve competitive results on any of our datasets. This shows that currently openly available GFMs are still far from being able to compete with classic GNNs on realistic datasets with rich node attributes.
@@ 5.3 High label rate, random and temporal data splits, transductive and inductive settings
First, the observations from the previous subsection about the usefulness of graph structure, benefits of attention-based GNNs, and weak performance of GFMs also apply to high label rate settings. Next, we observe that temporal data splits are significantly more challenging for all models than random ones (with the exception of the avazu-ctr dataset). This is important as in real-world applications temporal distributional shifts are common, and not considering them can provide overly optimistic performance estimates. Further, the considered models perform significantly worse in the inductive setting than in the transductive one. These observations highlight the importance of developing GML methods that are more resilient to temporal distributional shifts and dynamic changes in the graph structure for industrial applications. In the absence of such methods, frequent retraining of GNNs on new data is recommended to achieve the best results. Note that GFMs that only utilize in-context learning to adapt to new graphs do not suffer from the transductive/inductive mismatch and thus represent a promising direction, but their performance is currently not on par with GNNs on most datasets.
Overall, our main findings are the following:
  • GNNs can provide great benefits in industrial applications where graph information is available, with attention-based GNNs showing particularly strong results. However, GBDT models popular in industrial applications can serve as strong baselines when provided with graph information as additional input features, especially in regression tasks.
@@
  • The currently available general-purpose graph foundation models provide very inconsistent performance, occasionally achieving great results, but more often failing to compete with classic methods. Thus, the problem of creating truly general GFMs is far from being solved yet.
@@ 6 Conclusion
We introduce GraphLand — a set of diverse graph datasets representing realistic industrial applications of GML. Besides serving as an extended testbed for GNNs and GFMs, GraphLand allows investigating such previously underexplored research questions in the GML literature as how temporal distributional shifts and the inductive prediction setting influence the performance of GML methods. We hope GraphLand will encourage the evaluation of GML methods under more realistic and diverse settings, the development of GML methods that are more resilient to temporal distributional shifts and dynamically changing graph structure, and the development of more consistently performing GFMs that can handle different node feature sets that go beyond just textual descriptions.
@@ Acknowledgments
We thank Islam Umarov for providing the data for city-reviews dataset; Daniil Fedulov and Daniil Bondarenko for collecting the data for artnet-views and artnet-exp datasets; Dmitry Bikulov and Alexander Ulyanov for sharing the data for web-traffic, web-fraud, and web-topics datasets; Alexandr Ruchkin and Aleksei Istomin for collecting the data for city-roads-M and city-roads-L datasets.
@@ A.1 Dataset descriptions
In this section, we provide a more detailed description of the GraphLand datasets. Note that none of the proposed datasets contain any personal information. Our source code and instructions how to download our datasets can be found in our GitHub repository.
@@
web-traffic, web-fraud and web-topics   These three datasets are web-graphs — they represent a segment of the Internet. The nodes are websites, and a directed edge connects two nodes if at least one user followed a link from one website to the other in a selected period of time. We prepared three datasets with the same graph but different tasks: in web-traffic, the task is to predict how many users visited a website in a specific period of time (regression); in web-fraud, the task is to predict which websites are fraudulent (strongly imbalanced binary classification); and in web-topics, the task is to predict the topic that a website belongs to (multiclass classification). With almost $3$ million nodes, this is one of the largest publicly available attributed graphs that is not a citation network. Nodes in this graph have more than two hundred features, examples of which include the number of videos on the website (numerical feature), the website’s zone and whether the website is on a free hosting (categorical features).
@@
artnet-views and artnet-exp   These two datasets represent a social network of art creators. The nodes are users, and a directed edge connects two nodes if one user is subscribed to the other. We prepared two datasets with the same graph but different tasks: in artnet-views, the task is to predict how many views a user receives in a specific period of time (regression); and in artnet-exp, the task is to predict which users create explicit art content (binary classification). The examples of node features in this graph include user interests (categorical features).
@@
city-roads-M and city-roads-L   These datasets are obtained from the logs of a navigation service and represent the road networks of two major cities, with the second one being several times larger than the first. The nodes are segments of roads, and a directed edge connects two nodes if the segments are incident to each other and moving from one segment to the other is permitted by traffic rules. The task is to predict the average travel speed on a road segment at a specific timestamp (regression). The features include various information about the road segment such as binary indicators of whether there is a bike dismount sign, whether the road segment ends with a crosswalk or a toll post, whether the road segment is in poor condition, whether it is restricted for trucks, and whether it has a mass transit lane (categorical features). The examples of numerical features are the length of the road and the geographic coordinates of the road endpoints.
@@
city-reviews   This dataset is obtained from the logs of a review service in which users can leave reviews and ratings for places and organizations in two major cities. The nodes are users and an edge connects two nodes if the users often leave reviews for the same organizations. The graph is undirected. The task is fraud detection — to predict which users leave fraudulent reviews (binary classification). The node features are based on user interactions with the service and their examples include the share of negative reviews among all reviews the user has left (numerical feature) and the browser that is used to access the service by the user (categorical feature).
@@
avazu-ctr   This dataset is based on open data that has been introduced at the Kaggle competition organized by Avazu (Wang & Cukierski, 2014). The data contains information about interactions between devices used to access the Internet, websites, and advertisements. In our graph, the nodes are devices, and an edge connects two nodes if the devices often visit the same websites. The graph is undirected. A smaller version of a similar dataset has been used by Ivanov & Prokhorenkova (2021); however, it contained only a small subset of devices, while for our dataset we collected data for all the available devices which makes our graph more than 50 times larger. The task is to predict the advertisement click-through rate (CTR) observed on devices (regression). Nodes in this graph have more than two hundred numerical features; however, most of them were anonymized in the original data source.
@@
hm-categories and hm-prices   These datasets are based on open data that has been introduced at the Kaggle competition organized by H&M (Garc´ıa Ling et al., 2022). The graph represents a co-purchasing network. The nodes are products, and an edge connects two nodes if the products are often bought by the same customers. The graph is undirected. We prepared two datasets with the same graph but different tasks: in hm-categories, the task is to predict the product category (multiclass classification), and in hm-prices, the task is to predict the product price (regression). The node features in this dataset include product metadata such as product color (categorical feature), as well as information obtained from product purchasing statistics such as what proportion of product purchases occurs on different weekdays (numerical features).
@@
pokec-regions   This dataset is based on the data from Takac & Zabovsky (2012). It represents the online social network Pokec. The nodes are users, and a directed edge connects two nodes if one user has marked the other one as a friend. While this graph is quite popular in network analysis as an example of a classic social network, it is relatively rarely used for machine learning, with the exception of Lim et al. (2021) who use the same graph as us but with different task, node features, and data split. In our dataset, the task is to predict which region a user is from (extreme multiclass classification with $183$ classes). The node features in our dataset are based on user profile information, examples of them include the profile completion proportion (numerical feature) and binary indicators of whether different profile fields are filled (categorical features).
@@
twitch-views   This dataset is based on the data from Rozemberczki & Sarkar (2021). It represents the live-streaming network Twitch. The nodes are users, and a directed edge connects two nodes if one user is subscribed to the other. The task is to predict how many views a user gets in a specific period of time (regression). The node features are based on user profile information and examples of them include user language and affiliate status (categorical features).
@@
tolokers-2   This is a new version of the dataset tolokers from Platonov et al. (2023b); Likhobaba et al. (2023) with a significantly extended set of node features. It is based on the data from the Toloka crowdsourcing platform. The nodes are tolokers, and an edge connects two nodes if these tolokers have worked on the same task. The graph is undirected. The task is fraud detection — to predict which tolokers have been banned in one of the projects (binary classification). The new node features include various performance statistics of workers, such as the number of approved assignments and the number of skipped assignments (numerical features), as well as worker’s profile information, such as their education level (categorical feature).
@@ A.2 Dataset properties
A key characteristic of our benchmark is its diversity. As described above, our graphs come from different domains and have different prediction tasks. Their edges are also constructed in different ways (based on user interactions, activity similarity, physical connections, etc.). However, the proposed datasets also differ in many other ways. Some properties of our graphs are presented in Table 3 (see below for the details on how the provided characteristics are defined). First, note that the sizes of our datasets range from 1111K to 33M nodes. The smaller graphs can be suitable for compute-intensive models, while the larger graphs can provide a moderate scaling challenge. The average and median degrees of our graphs also vary significantly and our benchmark has both sparse and relatively dense graphs, including graphs with the average degree in the order of hundreds which is larger than the average degrees of most datasets used in current GML research (such graphs may highlight the importance of attention-based GNNs with their soft edge selection mechanisms). The average distance between two nodes in our graphs varies from $2.45$ for hm-categories and hm-prices to $194$ for city-roads-L; and graph diameter (maximum distance) varies from $8$ for twitch-views to $553$ for city-roads-L. Further, we report the values of clustering coefficients which show how typical closed node triplets are for the graph. In the literature, there are two definitions of clustering coefficients (Boccaletti et al., 2014): the global clustering coefficient and the average local clustering coefficient. We have both graphs where the clustering coefficients are high and graphs where they are almost zero, as well as graphs where global and local clustering coefficients significantly disagree (which is possible for graphs with imbalanced degree distributions). The degree assortativity coefficient is defined as the Pearson correlation coefficient of degrees among pairs of linked nodes. For most of our graphs, the degree assortativity is either negative or close to zero, which means that nodes do not tend to connect to other nodes with similar degrees, while city-roads-M and city-roads-L datasets are the exceptions — for them the degree assortativity is positive and large.
@@
Further, let us discuss the graph-label relationships in our datasets. To measure the similarity of labels of connected nodes for regression datasets, we use target assortativity — the Pearson correlation coefficient of target values between pairs of connected nodes. For instance, for the city-roads-M and city-roads-L datasets, the target assortativity is positive and quite large, which shows that nodes tend to connect to other nodes with similar target values (which is expected for the task of speed prediction in road networks), while for the twitch-views and web-traffic datasets, the target assortativity is negative. For classification datasets, the similarity of neighbors’ labels is usually called homophily: in homophilous datasets, nodes tend to connect to nodes of the same class. How to properly measure homophily has recently attracted some research. It has been noted by Lim et al. (2021) and Platonov et al. (2023a) that homophily measures typically used in the literature — such as the proportion of edges connecting nodes of the same class — are not appropriate for comparing homophily levels between graphs with different numbers of classes and their size balance. Platonov et al. (2023a) proposed a set of properties that a homophily measure appropriate for use in such comparisons should satisfy and Mironov & Prokhorenkova (2024) constructed the first known homophily measure that satisfies all these properties — unbiased homophily. Thus, in our work, we use unbiased homophily to measure the homophily levels of our datasets. Unbiased homophily (with $\alpha=0$, see Mironov & Prokhorenkova (2024) for more details) takes values in [1, 1][-1,\;1] with $1$ indicating perfect homophily, 1-1 indicating perfect heterophily, and 0 indicating no preference between homophilous and heterophilous edges (such graphs are typically referred to as heterophilous in the literature, although a more appropriate term would be non-homophilous). Note that the values of unbiased homophily should not be compared to values of other homophily measures used in the literature; the unbiased homophily levels for some popular graph node classification datasets are provided in Mironov & Prokhorenkova (2024). Unbiased homophily indicates that among our datasets pokec-regions and city-reviews are homophilous, while the other ones are non-homophilous. Thus, our benchmark significantly expands the set of available non-homophilous graph datasets.
@@
Table 3: Characteristics of the proposed GraphLand datasets.
node classification node regression
hm-categories
pokec-regions
web-topics
tolokers-2
city-reviews
artnet-exp
web-fraud
hm-prices
avazu-ctr
city-roads-M
city-roads-L
twitch-views
artnet-views
web-traffic
median degree $45$ $13$ 22 $30$ 44 22 22 $45$ $71$ 44 33 $32$ 22 22
diameter $13$ $14$ $36$ $11$ 1919 $13$ $36$ $13$ $14$ $383$ $553$ 88 $13$ $36$
# classes 2121 $183$ $28$ 22 22 22 22 N/A N/A N/A N/A N/A N/A N/A
# node features $35$ $56$ $263$ $16$ $37$ $75$ $266$ $41$ $260$ $26$ $26$ 44 $50$ $267$
@@
Average degree and median degree are the average and median numbers of neighbors a node has, respectively. Since all our graphs are connected (when treated as undirected graphs), for any two nodes there is a path between them. Average distance is the average length of the shortest paths between all pairs of nodes, while diameter is the maximum length of the shortest paths between all pairs of nodes. For our largest graphs — the ones used for the pokec-regions, web-traffic, web-fraud, and web-topics datasets — we approximate average distance with an average over distances for 100100K randomly sampled node pairs. Global clustering coefficient is computed as the tripled number of triangles divided by the number of pairs of adjacent edges (i.e., it is the fraction of closed triplets of nodes among all connected triplets). Average local clustering coefficient first computes the local clustering of each node, which is the fraction of connected pairs of its neighbors, and then averages the obtained values among all nodes. Degree assortativity is the Pearson correlation coefficient between the degrees of connected nodes. Further, target assortativity for regression datasets is the Pearson correlation coefficient between target values of connected nodes. For computing unbiased homophily, we follow Mironov & Prokhorenkova (2024) and use the simplest version of this measure with the $\alpha$ parameter set to 0.
Appendix B Experimental setting

### v4

@@ 5.2 Low label rate, random data split, transductive setting
RH $\makebox[20.00003pt][r]{$15.48$}\pm 1.93$ MLE MLE $\makebox[20.00003pt][r]{$31.83$}\pm 2.55$ $\makebox[20.00003pt][r]{$13.38$}\pm 2.59$ MLE
TS-GNN (ICL) TH $\makebox[20.00003pt][r]{$18.91$}\pm 0.32$ MLE MLE $\makebox[20.00003pt][r]{$12.59$}\pm 1.97$ $\makebox[20.00003pt][r]{$9.46$}\pm 1.25$ MLE
THI $\makebox[20.00003pt][r]{$18.91$}\pm 0.32$ MLE MLE $\makebox[20.00003pt][r]{$12.59$}\pm 1.97$ $\makebox[20.00003pt][r]{$9.46$}\pm 1.25$ MLE
@@
The results for experiments under the random low label rate data split (RL) are provided in Table 2.
@@
First, we can see that the graph structure in our datasets is very beneficial for the considered tasks as all GNNs always significantly outperform ResMLP and NFA-augmented models always significantly outperform their counterparts without NFA. Providing graph-agnostic models with graph information through NFA significantly improves their performance and even allows LightGBM-NFA to achieve the best results on four datasets including three regression ones. At the same time, GNNs outperform ResMLP-NFA on these datasets, which suggests that the success of LightGBM-NFA is due to GBDTs being better at handling numerical features and regression targets (Gorishniy et al., 2021, 2022; Zhang et al., 2023). This highlights that GBDTs with augmented features are strong baselines in realistic industrial settings. Still, GNNs manage to outperform them on most datasets, showing their effectiveness for industrial applications.
@@
As for GFMs, we find that all the considered models produce very weak results on all our datasets. This shows that currently available GFMs are still far from being able to compete with classic GNNs on realistic datasets with rich node attributes.
@@ 5.3 High label rate, random and temporal data splits, transductive and inductive settings
First, the observations from the previous subsection about the usefulness of graph structure, benefits of attention-based GNNs, and weak performance of GFMs also apply to high label rate settings. Next, we observe that temporal data splits are significantly more challenging for all models than random ones (with the exception of the avazu-ctr dataset). This is important as in real-world applications temporal distributional shifts are common, and not considering them can provide overly optimistic performance estimates. Further, the considered models perform significantly worse in the inductive setting than in the transductive one. These observations highlight the importance of developing GML methods that are more resilient to temporal distributional shifts and dynamic changes in the graph structure for industrial applications. In the absence of such methods, frequent retraining of GNNs on new data is recommended to achieve the best results. Note that GFMs that only utilize in-context learning to adapt to new graphs do not suffer from the transductive/inductive mismatch and thus represent a promising direction, but their performance is currently very weak compared with GNNs on all datasets.
Overall, our main findings are as follows:
  • GNNs can provide substantial benefits in industrial applications where graph information is available, with attention-based GNNs showing particularly strong results. However, GBDT models popular in industrial applications can serve as strong baselines when provided with graph information as additional input features, especially in regression tasks.
@@
  • The currently available general-purpose graph foundation models achieve very weak results on our datasets. Thus, the problem of creating truly general GFMs is far from being solved yet. We hope our benchmark can serve as a reliable testbed for future research in this direction, and first such successful use cases have already appeared (Eremeev et al., 2025a, b).
@@ 6 Conclusion
We introduce GraphLand — a set of diverse graph datasets representing realistic industrial applications of GML. Besides serving as an extended testbed for GNNs and GFMs, GraphLand allows investigating such previously underexplored research questions in the GML literature as how temporal distributional shifts and the inductive prediction setting influence the performance of GML methods. We hope GraphLand will encourage the evaluation of GML methods under more realistic and diverse settings, the development of GML methods that are more resilient to temporal distributional shifts and dynamically changing graph structure, and the development of better performing GFMs that can handle different node feature sets that go beyond textual descriptions.
@@ Acknowledgments
We thank Islam Umarov for providing the data for city-reviews dataset; Daniil Fedulov and Daniil Bondarenko for collecting the data for artnet-views and artnet-exp datasets; Dmitry Bikulov and Alexander Ulyanov for sharing the data for web-traffic, web-fraud, and web-topics datasets; Alexandr Ruchkin and Aleksei Istomin for collecting the data for city-roads-M and city-roads-L datasets; Fedor Velikonivtsev for helping with data processing for city-roads-M and city-roads-L datasets. We also thank Dmitry Eremeev for useful discussions.
@@ A.1 Dataset description
In this section, we provide a more detailed description of the GraphLand datasets. Note that none of the proposed datasets contain any personal information. For the datasets based on data newly collected for our benchmark, the public release of the data was approved by a legal team. GraphLand datasets are available at Zenodo. Our source code and instructions on how to reproduce our experimental results are available in our GitHub repository.
@@
web-fraud, web-topics, and web-traffic   These three datasets are web-graphs — they represent a segment of the Internet. The nodes are websites, and a directed edge connects two nodes if at least one user followed a link from one website to the other in a selected period of time. We prepared three datasets with the same graph but different tasks: in web-fraud, the task is to predict which websites are fraudulent (strongly imbalanced binary classification); in web-topics, the task is to predict the topic that a website belongs to (multiclass classification); and in web-traffic, the task is to predict how many users visited a website in a specific period of time (regression). With almost $3$ million nodes, this is one of the largest publicly available attributed graphs that is not a citation network. Nodes in this graph have more than two hundred features, examples of which include the number of videos on the website (numerical feature), the website’s zone, and whether the website is on a free hosting (categorical features). Data for these datasets was obtained from the Yandex search engine.
@@
artnet-views and artnet-exp   These two datasets represent a social network of art creators. The nodes are users, and an undirected edge connects two nodes if the users are friends. We prepared two datasets with the same graph but different tasks: in artnet-views, the task is to predict how many views a user receives in a specific period of time (regression); and in artnet-exp, the task is to predict which users create explicit art content (binary classification). Examples of node features include user interests (categorical features). Data for these datasets was obtained from the generative AI art creation platform Shedevrum.
@@
city-roads-M and city-roads-L   These datasets are obtained from the logs of a navigation service and represent the road networks of two major cities, with the second one being several times larger than the first. The nodes are segments of roads, and a directed edge connects two nodes if the segments are incident to each other and moving from one segment to the other is permitted by traffic rules. The task is to predict the average travel speed on a road segment at a specific timestamp (regression). The features include various information about the road segment such as binary indicators of whether there is a bike dismount sign, whether the road segment ends with a crosswalk or a toll post, whether the road segment is in poor condition, whether it is restricted for trucks, and whether it has a mass transit lane (categorical features). Numerical features include the length of the road segment and the geographic coordinates of the road endpoints. Data for these datasets was obtained from the Yandex Maps service.
@@
city-reviews   This dataset is obtained from the logs of a review service in which users can leave reviews and ratings for places and organizations in two major cities. The nodes are users, and an undirected edge connects two nodes if the users often leave reviews for the same organizations. The task is fraud detection — to predict which users leave fraudulent reviews (binary classification). The node features are based on user interactions with the service and their examples include the share of negative reviews among all reviews left by a user (numerical feature) and the browser that is used to access the service by the user (categorical feature). Data for this dataset was obtained from the Yandex Maps service.
@@
avazu-ctr   This dataset is based on open data that has been introduced at the Kaggle competition organized by Avazu (Wang & Cukierski, 2014). The data contains information about interactions between devices used to access the Internet, websites, and advertisements. In our graph, the nodes are devices, and an edge connects two nodes if the devices often visit the same websites. The graph is undirected. A smaller version of a similar dataset has been used by Ivanov & Prokhorenkova (2021); however, it contained only a small subset of devices, while for our dataset we collected data for all the available devices, making our graph more than 50 times larger. The task is to predict the advertisement click-through rate (CTR) observed on devices (regression). Nodes in this graph have more than two hundred numerical features; however, most of them were anonymized in the original data source.
@@
hm-categories and hm-prices   These datasets are based on open data that has been introduced at the Kaggle competition organized by H&M (García Ling et al., 2022). The graph represents a co-purchasing network. The nodes are products, and an edge connects two nodes if the products are often bought by the same customers. The graph is undirected. We prepared two datasets with the same graph but different tasks: in hm-categories, the task is to predict the product category (multiclass classification), and in hm-prices, the task is to predict the product price (regression). The node features in this dataset include product metadata such as product color (categorical feature), as well as information obtained from product purchasing statistics such as what proportion of product purchases occurs on different weekdays (numerical features).
@@
pokec-regions   This dataset is based on the data from Takac & Zabovsky (2012). It represents the online social network Pokec. The nodes are users, and a directed edge connects two nodes if one user has marked the other one as a friend. While this graph is quite popular in network analysis as an example of a classic social network, it is relatively rarely used for machine learning, with the exception of Lim et al. (2021) who use the same graph as we do but with different task, node features, and data split. In our dataset, the task is to predict which region a user is from (extreme multiclass classification with $183$ classes). The node features in our dataset are based on user profile information, examples of them include the profile completion proportion (numerical feature) and binary indicators of whether different profile fields are filled (categorical features).
@@
twitch-views   This dataset is based on the data from Rozemberczki & Sarkar (2021). It represents the live-streaming network Twitch. The nodes are users, and an undirected edge connects two nodes if both users follow each other. The task is to predict how many views a user receives in a specific period of time (regression). The node features are based on user profile information and examples of them include user language and affiliate status (categorical features).
@@
tolokers-2   This is a new version of the dataset tolokers from Platonov et al. (2023b); Likhobaba et al. (2023) with a significantly extended set of node features. It is based on the data from the Toloka crowdsourcing platform and the graph represents a network of tolokers (workers). The nodes are tolokers, and an edge connects two nodes if these tolokers have worked on the same task. The graph is undirected. The task is fraud detection — to predict which tolokers have been banned in one of the projects (binary classification). The new node features include various performance statistics of workers, such as the number of approved assignments and the number of skipped assignments (numerical features), as well as workers’ profile information, such as their education level (categorical feature).
Some graphs in our benchmark are undirected, and some are directed (see individual dataset descriptions above). All our undirected graphs consist of a single connected component, and all our directed graphs consist of a single weakly connected component.
@@ A.2 Dataset properties
A key characteristic of our benchmark is its diversity. As described above, our graphs come from different domains and have different prediction tasks. Their edges are also constructed in different ways (based on user interactions, activity similarity, physical connections, etc.). However, the proposed datasets also differ in many other ways. Some properties of our graphs are presented in Table 1 (see below for the details on how the provided characteristics are defined). First, note that the sizes of our datasets range from 1111K to 33M nodes. The smaller graphs can be suitable for compute-intensive models, while the larger graphs can provide a moderate scaling challenge. The average and median degrees of our graphs also vary significantly and our benchmark has both sparse and relatively dense graphs, including graphs with the average degree in the order of hundreds which is larger than the average degrees of most datasets used in current GML research (such graphs may highlight the importance of attention-based GNNs with their soft edge selection mechanisms). The average distance between two nodes in our graphs varies from $2.45$ for hm-categories and hm-prices to $194$ for city-roads-L; and graph diameter (maximum distance) varies from $8$ for twitch-views to $553$ for city-roads-L. Further, we report the values of clustering coefficients, which show how typical closed node triplets are for the graph. In the literature, there are two definitions of clustering coefficients (Boccaletti et al., 2014): the global clustering coefficient and the average local clustering coefficient. We have both graphs where the clustering coefficients are high and graphs where they are almost zero, as well as graphs where global and local clustering coefficients significantly disagree (which is possible for graphs with imbalanced degree distributions). The degree assortativity coefficient is defined as the Pearson correlation coefficient of degrees among pairs of linked nodes. For most of our graphs, the degree assortativity is either negative or close to zero, which means that nodes do not tend to connect to other nodes with similar degrees, while city-roads-M and city-roads-L datasets are the exceptions — for them the degree assortativity is positive and large.
@@
Further, let us discuss the graph-label relationships in our datasets. To measure the similarity of labels of connected nodes for regression datasets, we use target assortativity — the Pearson correlation coefficient of target values between pairs of connected nodes. For instance, for the city-roads-M and city-roads-L datasets, the target assortativity is positive and quite large, which shows that nodes tend to connect to other nodes with similar target values (which is expected for the task of speed prediction in road networks), while for the twitch-views and web-traffic datasets, the target assortativity is negative. For classification datasets, the similarity of neighbors’ labels is usually called homophily: in homophilous datasets, nodes tend to connect to nodes of the same class. How to properly measure homophily has recently attracted some research. It has been noted by Lim et al. (2021) and Platonov et al. (2023a) that homophily measures typically used in the literature — such as the proportion of edges connecting nodes of the same class — are not appropriate for comparing homophily levels between graphs with different numbers of classes and their size balance. Platonov et al. (2023a) proposed a set of properties that a homophily measure appropriate for use in such comparisons should satisfy and Mironov & Prokhorenkova (2024) constructed the first known homophily measure that satisfies all these properties — unbiased homophily. Thus, in our work, we use unbiased homophily to measure the homophily levels of our datasets. Unbiased homophily (with $\alpha=0$, see Mironov & Prokhorenkova (2024) for more details) takes values in [1, 1][-1,\;1] with $1$ indicating perfect homophily, 1-1 indicating perfect heterophily, and 0 indicating no preference between homophilous and heterophilous edges (such graphs are typically referred to as heterophilous in the literature, although a more appropriate term would be non-homophilous). Note that the values of unbiased homophily should not be compared to values of other homophily measures used in the literature; the unbiased homophily levels for some popular graph node classification datasets are provided by Mironov & Prokhorenkova (2024). Unbiased homophily indicates that among our datasets pokec-regions and city-reviews are homophilous, while the remaining datasets are non-homophilous. Thus, our benchmark significantly expands the set of available non-homophilous graph datasets.
@@
Average degree and median degree are the average and median numbers of neighbors a node has, respectively. Since all our graphs are connected (when treated as undirected graphs), for any two nodes there is a path between them. Average distance is the average length of the shortest paths between all pairs of nodes, while diameter is the maximum length of the shortest paths between all pairs of nodes. For our largest graphs — the ones used for the pokec-regions, web-traffic, web-fraud, and web-topics datasets — we approximate the average distance with the average over distances for 100100K randomly sampled node pairs. Global clustering coefficient is computed as three times the number of triangles divided by the number of pairs of adjacent edges (i.e., it is the fraction of closed triplets of nodes among all connected triplets). Average local clustering coefficient first computes the local clustering of each node, which is the fraction of connected pairs of its neighbors, and then averages the obtained values among all nodes. Degree assortativity is the Pearson correlation coefficient between the degrees of connected nodes. Further, target assortativity for regression datasets is the Pearson correlation coefficient between target values of connected nodes. For computing unbiased homophily, we follow Mironov & Prokhorenkova (2024) and use the simplest version of this measure with the $\alpha$ parameter set to 0.
Appendix B Experimental setup